OpenStudy (abbycross167):

Can someone please help me with this question? I will fan and give a medal... The width, w, of a rectangular rug is 4 less than its length, L. If the area of the rug is 120 square feet, which equation could be used to find the dimensions of the rug? a. L(4-L) = 120 b. L(L-4) = 120 c. 2(L-4) + 2L = 120 d. 2w + 2L = 120

OpenStudy (isaidavila):

The área of this rectangular rug is A=b.h or A=w.L So If the lenght is L Then width is L-4 So use that to solve that

OpenStudy (abbycross167):

I'm not sure how to...

OpenStudy (isaidavila):

Ok A=L(L-4) 120=L(L-4)

OpenStudy (abbycross167):

@isaidavila I don't understand still... could you show me how the answer is a?

jhonyy9 (jhonyy9):

@abbycross167 do you understand how you can solve it ?

jhonyy9 (jhonyy9):

are you here please ?

OpenStudy (abbycross167):

@jhonyy9 I am sorry my wifi dosen't very good signal. But no sir/ma'am I'm not the best at figuring out equations

jhonyy9 (jhonyy9):

so how you can continue it ?

OpenStudy (abbycross167):

what do you mean?

jhonyy9 (jhonyy9):

from the text of your exercise you know that the width is leas by 4 than length so w=l-4 and you know that area of a rectangle is equal l*w = l(l-4) area you know equal120 ft^2

jhonyy9 (jhonyy9):

can you calculi from theses l=?

jhonyy9 (jhonyy9):

you know that 120=l(l-4) - yes ? how you get a quadratic ?

jhonyy9 (jhonyy9):

so what is the answer to your exercise ? because i see that you not need solve this quadratic just get the equation for what solved can calculi the length and the width of rectangle yes ?

OpenStudy (abbycross167):

oh ok, thank you for your help!!

jhonyy9 (jhonyy9):

np was my pleasure than these was on your help really true good luck bye